i have two rectangles that have the same area. one's width is x and its length is 2x+2 and the other's width is x+3 and its length is x+1. i need to find x
Answer by richwmiller(17219)

x*(2x+2)=(x+3)*(x+1)
2x^2+2x = x^2+4x+3
x^2-2x-3 = 0
x=-1 disregard
which leaves
x=3
(3,8) and (6,4)